OpenStudy (anonymous):

A tortoise is walking in the desert. It walks 9.6 meters in 6 minutes. What is its speed?

OpenStudy (chosenmatt):

what do you think so far?

OpenStudy (perl):

speed = total distance / total time

OpenStudy (chosenmatt):

what is tortoise doing on a desert in the first place? XD

OpenStudy (anonymous):

speed=9.6/6

OpenStudy (acxbox22):

we have the value for total distance and total time so we can plug them into the equation and solve for speed

OpenStudy (anonymous):

1.6

OpenStudy (perl):

correct, but make sure to label the units,

OpenStudy (acxbox22):

yes now divide 9.6/6 to get the speed in meters per minute

OpenStudy (perl):

1,6 what?

OpenStudy (chosenmatt):

^^^

OpenStudy (anonymous):

1.6 meters per min

OpenStudy (perl):

yes :)

OpenStudy (perl):

1.6 meter/min

OpenStudy (anonymous):

thanks i have another one can you help me with

OpenStudy (perl):

you can convert to mph , if you are american that is more familiar

OpenStudy (perl):

ok

OpenStudy (anonymous):

do you want me to open a new page

OpenStudy (perl):

not necessary

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

As Jupiter revolves around the sun, it travels at a speed of approximately 8 miles per second. Convert this speed to miles per minute. At this speed, how many miles will Jupiter travel in 4 minutes? Do not round your answers. find speed distance traveled in 4 min

OpenStudy (perl):

we can use google to convert type the following into the search box: 8 mile/sec = mile/min

OpenStudy (anonymous):

480

OpenStudy (perl):

alternatively you can multiply by conversion factor 8 mile / sec * 60 sec / 1 min = 480 mile / min

OpenStudy (perl):

correct

OpenStudy (anonymous):

so thats speed

OpenStudy (anonymous):

?

OpenStudy (perl):

right

OpenStudy (perl):

just different units for the same speed

OpenStudy (perl):

At this speed, how many miles will Jupiter travel in 4 minutes? Do not round your answers. multiply the speed by the time 480 mile/min * 4 min = ? note that the units cancel , which is a nice check you should get an answer in miles

OpenStudy (anonymous):

120

OpenStudy (perl):

480 * 4 = 1920

OpenStudy (perl):

let me use parentheses (480 mile/min) * 4 min

OpenStudy (anonymous):

i / it woops

OpenStudy (anonymous):

so speed is 480 mi is 1920

OpenStudy (perl):

correct